OpenStudy (anonymous):

Choose the correct statement below about the density of water. The density of water under standard conditions is 10 kilograms per meter cubed or 1 gram per centimeter cubed. The density of water under standard conditions is 1000 kilograms per meter cubed or 103 gram per centimeter cubed. The density of water under standard conditions is 1000 kilograms per meter cubed or 1 gram per centimeter cubed. The density of water under standard conditions is 1003 kilograms per meter cubed or 1 gram per centimeter cubed.

OpenStudy (anonymous):

\[1000 \left [ \rm kg \over m^3 \right ] \cdot \left [ 1000 \rm g \over kg \right] \cdot \left[1 \rm m^3 \over 100^3 cm^3 \right] = {1000^2 \over 1000^2} \left [ \rm g \over cm^3 \right ]\]

OpenStudy (anonymous):

the second one :)

OpenStudy (anonymous):

Really? It's the third one, clearly.

OpenStudy (anonymous):

ofcourse third one....1000kg/1m^3=(10^6)grams/(10^6)cm^3=1g/cm^3